# Knowledge Transfer from High-Resource to Low-Resource Programming Languages for Code LLMs

### v5

@@ 3.1. Further Training on Natural Data
In figure 2(b), we do another experiment with The Stack that lends itself to a direct comparison with MultiPL-T. We randomly sample data from The Stack to get approximately the same volume of data that we generate with MultiPL-T. Thus, fine-tuning on these datasets will use similar computing resources as fine-tuning a model with MultiPL-T data. We use this data to fine-tune three versions on StarCoderBase-1B for six epochs, and evaluate the models at each epoch. We still get poor results with The Stack: Lua and OCaml performance barely increases whereas Racket performance decreases. In contrast, fine-tuning with MultiPL-T (dashed lines) will show significant gains.
@@
    (foldr + 0 (map string->number (string-split (number->string num) ""))))
@@
      (map char->integer
        (string->list (number->string num))) 10)))
@@
(d) Uses a function from the prompt, despite separators.
@@
(e) Several errors, e.g., repeat is hallucinated and uses a function from prompt.
@@ 3.2. Self-Instruction for Low-Resource Programming Languages
An alternative to fine-tuning on natural data is to fine-tune on LLM-generated data (Wang et al., 2023; Luo et al., 2023). The usual approach is to seed a model with hand-selected programs, prompt the model to generate more examples, and iterate until a large enough dataset has been collected. This type of approach has been used successfully to generate training data for Code LLMs in high-resource languages (Luo et al., 2023). However, we argue that self-instruction and its variations are unlikely to work for low-resource languages because Code LLMs often generate low-quality code in low-resource languages.
To illustrate, we use StarCoderBase-15B to generate functions in Racket, mimicking the first step of self instruction. We prompt the model with five hand-written examples (Appendix A) and have it generate five more functions, shown in Figure 4. Unfortunately, we find that three of the five model-generated programs have bugs, which we explain in the captions. This is a much higher error rate than what is evident from self-instruct datasets for high-resource languages (Chaudhary, 2023; Muennighoff et al., 2023).
Self-instruction and training further on existing public data do not help Code LLMs perform better on low-resource programming languages. A final alternative is to train on proprietary data, which does not lend itself to the kind of detailed evaluation that we do (§§ 3.1 and 5). Thus we now turn to the MultiPL-T approach.
Figure 5. The MultiPL-T pipeline for generating semi-synthetic data. Starting with a tested Python function \⃝raisebox{-0.9pt}{1}, we translate the header and test cases into each target language \⃝raisebox{-0.9pt}{2}. We prompt the model with the translated header and comment to generate 50-100 candidate translations \⃝raisebox{-0.9pt}{3} (varies by language). We append the compiled test cases to the candidate translations \⃝raisebox{-0.9pt}{4} and evaluate with the target language runtime. We deduplicate all programs that pass all test cases \⃝raisebox{-0.9pt}{5} to build the fine-tuning dataset \⃝raisebox{-0.9pt}{6}.
@@ 4.1. Filtering Data from A High-Resource Language for Translation and Test Generation
Table 1. Size of the Python source dataset after each filtering step.
@@
    We exclude Python functions that do not have a docstring or use non-ASCII characters. One could generalize to include functions that have an associated comment. However, we still end up with up over 5M candidate functions with this simple filter.
@@
    We use the Pyright (Pyright, 2023) Python checker to validate that each function returns a value, uses only the Python standard library, and is thus likely type-correct [^6] Pyright is a heuristic type-checker that does not guarantee soundness.. This narrows the 5M functions to approximately 460,000 functions.
@@
(a) CDF of test suite sizes
(b) CDF of test suite coverage
Figure 6. Python test suite sizes and coverage distribution among the unfiltered test generation dataset
@@
    """Count the number of Xs in a word."""
@@
Figure 7. An example prompt to generate a unit test where the code and comment are inconsistent: the comment counts X’s, but the code counts Y’s. Using a Code LLM to generate tests helps expose the inconsistency: we get some tests that are based on the comment and others based on the code.
@@ 4.2. Generating Python Unit Tests
The next step in MultiPL-T is to generate unit tests for each Python function. We will then compile these unit tests to target low-resource languages using the test case translators from MultiPL-E. However, instead of using a traditional test generator that synthesizes tests from code (as in (Lukasczyk et al., 2020)), we use a Code LLM to generate test cases by prompting the model to produce an assertion. The Code LLM conditions on the source code text and thus serves as a weak detector for inconsistencies between code and comments. For example, figure 7 shows a function that would be a bad training item: the code counts Y’s, but has a comment that says it counts X’s. When we generate several test cases independently, we end up with tests for both X and Y, which lowers the ratio of passing tests. Conditioning on the text makes it less likely that inconsistent tests will be generated, decreasing the likelihood that functions will be filtered out of the training set due to low coverage (as described below).
We prompt StarCoder-15B to generate five independent test suites for each function at high temperature [^7] Temperature is a parameter that controls the randomness of the next-token predictions. A higher temperature results in more varied (less predictable) output, while a lower temperature produces more conservative (more predictable) results. (0.8) to get a diverse set of candidate tests. We parse each generated test suite and extract all test cases that are suitable for translation using MultiPL-E. We take the set of matching test cases and run each test in isolation in a container to verify that it passes, discarding any that fail. If no correct tests are generated, we discard the function. The result is a set of nearly 160,000 Python functions with at least one passing test case. Figure 5(a) shows the distribution of test suite sizes. The median number of test cases per function is 7, and the mean is 12.1.
@@
Given the dataset of Python functions with docstrings and test suites, our next step is to filter out functions with low test coverage. We use line coverage as the coverage metric, and exclude all functions with less than 90% line coverage. The result is a dataset of 133,668 functions with 90% line coverage from tests.
Since we start with nearly 160,000 functions, this implies that most of the generated test suites that work have high line coverage. In fact, most functions have 100% line coverage (Figure 5(b)). This stringent criterion ensures that the functions in our final set are not just correct but are also comprehensively tested, reinforcing the reliability of our dataset.
@@
The steps described above are sufficient to generate data for an untyped low-resource language (e.g., Racket or Lua). However, for a typed target (e.g., OCaml or Julia), we also need to infer types. [^8] Julia can be used with or without types, and our dataset has both typed and untyped examples. Even when the source language supports type inference, we need to infer Python types to translate test cases. For example, a Python function that consumes an optional integer can receive None or a number $n$. However, when we translate that function to OCaml, we must translate that number to Some($n$). We can only do this by inferring that the Python type is Optional[Int].
Our approach to type inference is simple: we deduce types based on test cases, ignoring the function body. We extract the instance type of each argument and expected return value in each test, computing the union type between the types at the same position among tests. For example, if the test cases apply foo(1) and foo(None), we infer Union[int, None] as the type of foo’s argument. Moreover, we simplify Union[T, None] to the more canonical Optional[T]. In the example above, Union[int, int, None] would be then simplified to Optional[int].
Following the steps above produces two datasets of Python functions—one with and one without type annotations—where every function has a docstring, and a suite of unit tests that achieve high coverage. These datasets can be used to generate training data for any low-resource language.
@@ 4.3. Translation from a High-Resource to a Low-Resource Language
    Docstring: We turn the Python docstring into a comment in the target language. The MultiPL-E toolchain translates between different comment formats, and also alters common type names in natural language using simple rules. For example, when translating from Python to OCaml, we turn “dictionary” into “association list”.
@@
    Function signature: We turn the Python function signature into a function signature in the target language. This involves translating types from Python into the target language and is trivial for untyped targets.
@@
    Original Python code: Finally, we add a comment (in the target language) that contains the original Python code. We find that this additional information increases the chance that the model generates a correct translation (§ 5.5).
@@
Given this prompt, we use StarCoderBase-15B to generate translations of each problem in our Python dataset. For all of our languages, we generate 50 translations [^9] For OCaml, we generated 100 translations per problem, as the base pass rate was significantly lower than other languages. at high temperature (0.8), to encourage the Code LLM to produce a more diverse set of candidate solutions (Chen et al., 2021).
@@ 4.4. Deduplication
We define a set of solutions to the same problem as diverse when they are distinct in form but behave the same on the test suite. Ensuring diversity in our generated and verified solutions is critical for ensuring that fine-tuned models learn a range of syntactic and semantic features in the target programming language. Moreover, redundant or similar solutions can dilute a dataset’s efficacy (Lee et al., 2022).
@@
Just resampling with high temperature does not guarantee diverse solutions: the LLM may still produce solutions that are nearly identical (e.g., with a few variables renamed). To address this, we employ a deduplication algorithm based on ROUGE-L (Lin and Och, 2004). ROUGE-L is a metric of text summarization quality, and quantifies the syntactic overlap between two pieces of text with a score ranging between 00 and $1$ where $1$ indicates highest similarity. We use a 0.6 as the similarity threshold for discarding duplicates. Before comparing a pair of solutions, we remove all comments from the code, as it may introduce noise in the deduplication process.
Running ROUGE-L on all pairs of items is prohibitively expensive. Instead, we use a heuristic that is amenable to parallelization (Algorithm 1). We apply ROUGE-L to deduplication items in fixed-size groups (we use size 200). Initially, the groups are solutions to the same prompt. We then randomly regroup items and run grouped deduplication again. The number of rounds of deduplication is proportional to the total number of items: more rounds increases the likelihood that duplicates will be removed. Ultimately, this results in a set of diverse, accurate, and semantically equivalent solutions for each prompt.
@@ 5. Evaluation
We use MultiPL-T data to fine-tune a large, medium, and small-sized Code LLM: CodeLlama-34B, StarCoderBase-15B, and StarCoderBase-1B. Both StarCoder models have the same architecture and were trained on exactly the same data. However, StarCoderBase-1B is significantly cheaper to train and evaluate, so we use it for the majority of the results in this section. In contrast, CodeLlama-34B is a larger model that was trained on an undisclosed dataset of publicly available code. We use it to demonstrate that our results generalize to larger Code LLMs.
@@
We fine-tune all models with a sequence length of 2,048 tokens. StarCoderBase-1B is fine-tuned for seven epochs with batch size $8$, with learning rate $3\times 10^{-5}$, $10$ steps warmup, and cosine learning rate decay. For StarCoderBase-15B and CodeLlama-34B, we make these configuration changes: ten epochs, batch size $32$, and learning rate $2\times 10^{-5}$.
@@
The work for this article was done over several months using V100 (32GB), A100 (80GB), and H100 (80GB) NVIDIA GPUs, as they were available across several clusters and servers. We estimate that we spent nearly 500 days of A100 (80GB) GPU time with the following breakdown:
  • Training: ~1,624 hours fine-tuning several versions of Code Llama-34B, StarCoderBase-15B, and StarCoderBase-1B. These include the models presenting in this section and in § 3.
  • Evaluation: ~230 hours running benchmarks, which include MultiPL-E and the new adversarial benchmark (§ 5.3).
  • Dataset Generation: ~9,984 hours generating the MultiPL-T training sets. This was the most significant use of resources, but is reusable for future model development.
@@
Language StarCoderBase-1B StarCoderBase-15B CodeLlama-34B
Base Fine-tuned Base Fine-tuned Base Fine-tuned
OCaml 1.5 9.7 6.9 19.9 18.3 27.4
Racket 4.7 11.3 11.8 21.0 15.9 29.1
R 5.4 8.9 10.2 17.3 18.2 25.5
Julia 11.3 15.6 21.1 35.2 31.8 43.5
Lua 12.1 17.3 26.6 31.0 38.1 43.9
Table 2. MultiPL-E pass@1 scores for 1B, 15B, and 34B parameter models before and after fine-tuned on MultiPL-T data. The fine-tuned models show significant improvement.
5.1. Fine-Tuning Small, Medium, and Large Code LLMs with MultiPL-T
Fine-tuning our base models with MultiPL-T-generated data on our target languages improves performance on MultiPL-E across the board (Table 2). For each model, we fine-tune a separate model for Julia, Lua, OCaml, R, and Racket. We checkpoint and evaluate the models at each epoch and report the peak performance. It is not a goal of this paper to maximize MultiPL-E scores. In fact, the next section suggests that it would be easy to improve some of these scores by either training longer on the data we already have or by letting MultiPL-T generate more data. Beyond the generally improved performance, we draw several other conclusions from our evaluation (Table 2):

### v6

@@ 3.1. Further Training on Natural Data
In figure 4(b), we do another experiment with The Stack that lends itself to a direct comparison with MultiPL-T. We randomly sample data from The Stack to get approximately the same volume of data that we generate with MultiPL-T. Thus, fine-tuning on these datasets will use similar computing resources as fine-tuning a model with MultiPL-T data. We use this data to fine-tune three versions on StarCoderBase-1B for six epochs, and evaluate the models at each epoch. We still get poor results with The Stack: Lua and OCaml performance barely increases and Racket performance decreases. In contrast, fine-tuning with MultiPL-T will show significant gains.
@@
(foldr +
0
(map string->number (string-split (number->string num) ""))))
@@
(map char->integer (string->list (number->string num))) 10)))
@@
(d) Utilizes what looks like a helper function, is-palindrome?, which is not defined.
@@
(e) The highlighted code calls two helper functions that are not defined.
@@ 3.2. Self-Instruction for Low-Resource Programming Languages
An alternative to fine-tuning on natural data is to fine-tune on LLM-generated data (Wang et al., 2023; Luo et al., 2024). The usual approach is to hand-select a seed dataset of programs, prompt the model with each seed to generate more programs, and iterate until a large enough dataset has been collected. This type of approach has been used successfully to generate training data for Code LLMs in high-resource languages (Luo et al., 2024). However, it should be obvious that these approaches presuppose that the LLM is good at generating reasonably correct and high-quality programs. We are interested in languages that the model is bad at, so it should not be surprising that self-instruction does not work.
Illustrating self-instruction
To illustrate how self-instruction goes wrong with low-resource languages, we use StarCoderBase-15B to generate functions in Racket, mimicking the first step of self-instruction. We prompt the model with five hand-written examples (included in the supplementary material) and have it generate five more functions (Figure 5). We find that four of the five model-generated programs have bugs. This is a much higher error rate than what is evident from self-instruct datasets for high-resource languages (Chaudhary, 2023; Muennighoff et al., 2024).
A self-instruction experiment
In the supplementary material we present results from an experiment where we self-instruct StarCoderBase-15B on Racket and get the expected poor results.
Self-instruction and training further on existing public data do not help Code LLMs perform better on low-resource programming languages. Thus we now turn to the MultiPL-T approach.
[…]
@@ 4.1. Filtering Data from a High-Resource Language for Translation and Test Generation
Table 1. Size of the Python source dataset after each filtering step.
@@
    We exclude Python functions that do not have a docstring or use non-ASCII characters. One could generalize to include functions that have an associated comment. However, we still end up with over 5M candidate functions with this simple filter.
@@
    We use the Pyright (Pyright, 2023) Python checker to validate that each function returns a value, uses only the Python standard library, and is thus likely type-correct. Pyright uses heuristics and makes no attempt at being sound. This does not impact MultiPL-T, since we merely use typeability as a heuristic for code quality. This narrows the 5M functions to approximately 460,000 functions.
@@
(a) The $y$-axis shows the fraction of test suites with fewer than $x$ tests.
(b) The $y$-axis shows the fraction of test suites with less than $x$coverage.
Figure 7. Python test suite sizes and coverage distribution among the unfiltered test generation dataset.
[…]
@@
"""Count the number of X’s in a word."""
@@
assert count("XXXXXX") == 6 # First completion
assert count("YYY") == 3 # Second completion
Figure 8. An example prompt to generate a unit test where the code and comment are inconsistent: the comment counts X’s, but the code counts Y’s. Using a Code LLM to generate tests helps expose the inconsistency: we get some tests that are based on the comment and others based on the code. We show two completions we got from StarCoderBase-15B in two successive queries. The “assert count(” is part of the prompt, but we show it with the completion for clarity.
[…]
@@ 4.2. Generating Python Unit Tests
The next step in MultiPL-T is to generate unit tests for each Python function. We will then compile these unit tests to target low-resource languages using the test case translators from MultiPL-E. We generate Python unit tests using the following steps.
  1. (1)
    We prompt the Code LLM to generate assertions given the Python function, taking care to generate several independent test suites to get high coverage, and executing the tests to ensure that they pass.
  2. (2)
    For each function, we measure the coverage of the aggregated test suites, discarding functions with less than 90% line coverage.
  3. (3)
    We use the tests to infer basic types for the Python function, which is necessary to compute the tests to certain target languages.
Test generation
Instead of using a traditional test generator that synthesizes tests from code (e.g., (Lukasczyk et al., 2020)), we use a Code LLM to generate test cases by prompting the model to produce an assertion. The Code LLM conditions on the source code text and thus serves as a weak detector for inconsistencies between code and comments. For example, figure 8 shows a function that would be a bad training item: the code counts Y’s, but has a comment that says it counts X’s. When we generate several test cases independently, we end up with tests for both X and Y, which lowers the ratio of passing tests. Conditioning on the text makes it less likely that inconsistent tests will be generated, decreasing the likelihood that functions will be filtered out of the training set due to low coverage (as described below).
We prompt StarCoder-15B to generate five independent test suites for each function with high temperature [^6] Temperature is a parameter that controls the randomness of the next-token predictions. A higher temperature results in more varied (less predictable) output, while a lower temperature produces more conservative (more predictable) results. (0.8) to get a diverse set of candidate tests. We parse each generated test suite and extract all test cases that are suitable for translation using MultiPL-E. We take the set of matching test cases and run each test in isolation in a container to verify that it passes, discarding any that fail. If no correct tests are generated, we discard the function. The result is a set of nearly 160,000 Python functions with at least one passing test case. Figure 7(a) shows the distribution of test suite sizes. The median number of test cases per function is 7, and the mean is 12.1. Note that we do not filter any tests after validated, so a function may be more tests that strictly necessary.
@@
Given the dataset of Python functions with docstrings and test suites, our next step is to filter out functions with low test coverage. We use line coverage as the coverage metric and exclude all functions with less than 90% line coverage. The result is a dataset of 133,668 functions with 90% line coverage from tests.
Since we start with nearly 160,000 functions, this implies that most of the generated test suites that work have high line coverage. In fact, most functions have 100% line coverage (Figure 7(b)). This stringent criterion ensures that the functions in our final set are not just correct but also comprehensively tested, reinforcing the reliability of our dataset. In the filtered Python dataset, the average function has 10.3 lines of code (SD=10.7) and on average 3.6 branches (SD=4.1).
def rep_or_hello(x):
"""Repeat string twice, or
produce Hello if no string given"""
if x is None:
return "Hello"
else:
return x + x
assert rep_or_hello(None) == "Hello"
assert rep_or_hello("y") = "yy"
(a) Python function and generated tests.
(* Repeat string twice, or produce
Hello if no string given. *)
let rep_or_hello (x : string option)
: string =
assert rep_or_hello None == "Hello";;
assert rep_or_hello (Some "y") = "yy"
(b) OCaml prompt and compiled tests.
Figure 9. An example that demonstrates the need for Python type inference when translating to OCaml. Figure 9(a) shows a Python function with generated test cases. We then compile both the Python function signature and the test cases to OCaml (figure 9(b)). However, to know that the Python string "y" must compile to Some "y" in OCaml, we need to infer Python types.
[…]
@@
The steps described above are sufficient to generate data for an untyped low-resource language (e.g., Racket or Lua). However, for a typed target (e.g., OCaml or Julia), we also need to infer types for two reasons. [^7] Julia can be used with or without types, and our dataset has both typed and untyped examples. Consider the case where we target OCaml. First, we rely on the LLM to generate an OCaml function body, given only a comment and the function header (let f x =). Without type annotations, the only signal about the desired type of f are the identifier names and the comment. With type annotations, the LLM is far more likely to produce a function with the expected type. Second, we need to infer Python types to compile test cases, which we illustrate in figure 9. In this example, we have a Python function that consumes an optional string. We have two test cases, one that applies the function to None and the other that applies the function to a string. When we compile these tests to OCaml, we have to transform the Python argument type Union[str,None] to the OCaml type string option. We can only do this type transformation after we have inferred the Python type.
Our approach to type inference is simple: we deduce types based on test cases, ignoring the function body. We extract the instance type of each argument and expected return value in each test, computing the union type between the types at the same position among tests. For example, if the test cases apply foo(1) and foo(None), we infer Union[int, None] as the type of foo’s argument. Moreover, we simplify Union[T, None] to the more canonical Optional[T]. For example, Union[int, int, None] would be then simplified to Optional[int]. This approach to type inference can only fail if the code is non-deterministic, which does not occur in our dataset.
Following the steps above produces two datasets of Python functions—one with and one without type annotations—where every function has a docstring and a suite of unit tests that achieve high coverage. These datasets can be used to generate training data for any low-resource language.
@@ 4.3. Translation from a High-Resource to a Low-Resource Language
    Docstring: We turn the Python docstring into a comment in the target language. The MultiPL-E toolchain translates between different comment formats and also alters common type names in natural language using simple rules. For example, when translating from Python to OCaml, we turn “dictionary” into “association list”.
@@
    Function signature: We turn the Python function signature into a function signature in the target language. This step may involve translating types from Python into the target language if they are required.
@@
    Original Python code: Finally, we add a comment (in the target language) that contains the original Python code. We find that this additional information increases the chance that the model generates a correct translation (§ 5.4.2).
@@
Given this prompt, we use StarCoderBase-15B to generate translations of each problem in our Python dataset. For all of our languages, we generate 50 translations [^8] For OCaml, we generated 100 translations per problem, as the base pass rate was significantly lower than other languages. with high temperature (0.8), to encourage the Code LLM to produce a more diverse set of candidate solutions (Chen et al., 2021).
@@ 4.4. Deduplication
We define a set of solutions as diverse when they differ in form. For instance, two solutions that perform identically on a test suite but differ in their implementation–one using a recursive function and the other a loop–are considered diverse. Ensuring diversity among generated and verified solutions is crucial for teaching fine-tuned models a variety of syntactic and semantic features of the target programming language. Furthermore, redundant or similar solutions may diminish the effectiveness of a dataset (Lee et al., 2022).
Just resampling with high temperature does not guarantee diverse solutions: the LLM may still produce nearly identical solutions (e.g., with a few variables renamed). To address this, we employ a deduplication algorithm based on ROUGE-L (Lin and Och, 2004). ROUGE-L is a metric of text summarization quality, and quantifies the syntactic overlap between two pieces of text with a score ranging between 00 and $1$ where $1$ indicates the highest similarity. We use 0.6 as the similarity threshold for discarding duplicates. Before comparing a pair of solutions, we remove all comments from the code, as it may introduce noise in the deduplication process.
Running ROUGE-L on all pairs of items is prohibitively expensive. Instead, we use a heuristic that is amenable to parallelization (Algorithm 1). We apply ROUGE-L to deduplication items in fixed-size groups (we use size 200). Initially, the groups are solutions to the same prompt, as this group is likely to have many duplicates. We then randomly regroup items and run grouped deduplication again. The number of rounds of deduplication is proportional to the total number of items: more rounds increase the likelihood that duplicates will be removed. Ultimately, this results in a set of diverse, accurate, and semantically equivalent solutions for each prompt.
@@ 5. Evaluation
In this section, we use MultiPL-T to fine-tune a variety of different models of varying sizes. We demonstrate state-of-the-art results on standard benchmarks for the natural language to code task (§ 5.2), novel qualitative and quantitative evaluation (§ 5.3), and ablations showing the significance of various design decisions in MultiPL-T(§ 5.4).
5.1. Experimental Setup and Implementation
@@
We fine-tune all models with a sequence length of 2,048 tokens. StarCoderBase-1B is fine-tuned for seven epochs with batch size $8$, with learning rate $3\times 10^{-5}$, $10$ steps warmup, and cosine learning rate decay. For StarCoderBase-15B, CodeLlama-34B, and CodeLlama-70B we make these configuration changes: ten epochs, batch size $32$, and learning rate $2\times 10^{-5}$.
@@
The work for this article was done over several months using V100 (32GB), A100 (80GB), and H100 (80GB) NVIDIA GPUs, as they were available across several clusters and servers. We estimate that we spent approximately 550 days of A100 (80GB) GPU time with the following breakdown:
  • Training: ~3,444 hours fine-tuning several versions of CodeLlama-70B, CodeLlama-34B, StarCoderBase-15B, and StarCoderBase-1B. These include the models presented in this section and in § 3.
  • Evaluation: ~310 hours running benchmarks, which include MultiPL-E and the new in-context learning benchmark (§ 5.3.1).
  • Dataset Generation: ~9,984 hours generating the MultiPL-T training sets. This was the most significant use of resources but is reusable for future model development. Using models larger than StarCoderBase-15B for generation would take up significantly more resources.
@@
5.2. Evaluation on Standard Benchmarks
Most Code LLM benchmarks target high resource programming languages, mostly Python. To evaluate our fine-tuned LLMs for low-resource languages, we use and extend MultiPL-E, which is the benchmark that was used to evaluate StarCoder, Code Llama, Stability.ai’s StableCode, and several other Code LLMs. As described in § 2.2, MultiPL-E is a benchmark for the natural language to code task. We report MultiPL-E results in the standard way: using the pass@1 metric, which is simply the mean pass rate on the benchmark, where a passing completion must successfully execute all tests.
Choosing baselines and comparisons
Before presenting results, we discuss how we believe fine-tuned models should be evaluated.
It is well known that larger models perform better because they can learn more complex patterns (Hoffmann et al., 2022). Moreover, there is a recent trend of building smaller models that outperform larger models by training them on far more data (de Vries, 2023). For example, StarCoder2-15B is the same size as StarCoderBase-15B, but is trained on 400% more data and performs 50% better on MultiPL-E (Lozhkov et al., 2024).
Therefore, it trivial to achieve an impressive fine-tuning result by simply starting with a newer, better model. We argue that the right way to evaluate a fine-tuning approach is to do all of the following: